# Electric conductivity of hot and dense nuclear matter

### v1

@@ 1 Introduction
Two approaches to studying hot and dense nuclear matter include the observation and description of electromagnetic probes, and the characterization of the medium with transport coefficients. Dileptons and photons are produced throughout the evolution of the nuclear fireball with little final state interaction, providing a probe of the electromagnetic spectral function ($\rho_{EM}$) in the medium [3]. The production of low-mass dilepton and photon radiation has been well described at multiple experiments [4, 5, 6]. Transport coefficients detail the long-wavelength properties of the medium, describing the flow of conserved charges throughout the collision. In this paper, we calculate the electric conductivity ($\sigma_{el}$), which can be extracted from the zero-energy limit of $\rho_{EM}$ at vanishing 3-momentum [7, 8, 9]. A close relationship exist between electromagnetic probes and the conductivity, as both are proportional to $\rho_{EM}$. In fact, future plans to measure very-low mass and momentum dileptons at the Schwer-Ionen Synchrotron (SIS), RHIC, and the LHC have potential to probe the conductivity’s transport peak [10, 11, 12]. These initiatives encourage the study of the conductivity, especially considering the significant variation in calculations using different methodologies [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@ 4 Gauge invariance in medium
Within the VDM, the $\rho$ meson couples to a conserved current, requiring it to be 4-dimensionally transverse, $q_{\mu}\Sigma_{\rho}^{\mu u}=0$. The Ward identities (Eq. 17) must be satisfied by the $\rho\pi\pi$ ($\Gamma_{\mu\,ab3}^{(3)}$) and $\rho\rho\pi\pi$ ($\Gamma_{\mu\nu\,ab33}^{(4)}$) vertices to ensure transversality:
@@
We start by assuming the corrections can be written as a sum of terms proportional to $\Sigma_{\pi}$. For the spacial components of the 3-point corrections, we additionally assume each term is proportional to a momentum $k_{i}$ or $q_{i}+k_{i}$, and can be written as a function of $k$ or $q+k$, as was the case in Ref. [26]. For the 3-point vertex we arrive a the corrections:
@@
where, $R=\rho,\sigma,K$. Here, we have divided each term in Eq. 22 by either $k_{0}|\vec{k}|$ or $(q_{0}+k_{0})|\vec{q}+\vec{k}|$, so that the corrections will have units of energy, and each term will be a function of either $k$ or $q+k$. This choice assures the corrections are finite in the limit where $k_{0}$ or $|\vec{k}|$ approach zero, without introducing additional parameters. These restrictions combined with Eq. 19 determine effective 3-point corrections. For the 4-point corrections we assume $\Gamma_{ij\,ab}^{\prime(4\,R\pi)}=0$ and use Eq. 20 to fix $\Gamma_{i0\,ab}^{\prime(4\,\pi)}$ and $\Gamma_{00\,ab}^{\prime(4\,\pi)}$, giving:
@@ 5 EM spectral function at finite temperature and density
Finally, we extract the conductivity using Eq. 9. In the right panel of figure 4 we plot the conductivity over temperature along the phase transition line. We provide results with and without vertex corrections, which provide a $\sim 15\%$ increase. This is because the corrections introduce additional channels that the EM current can couple to the medium through [25]. Although, this result is somewhat larger than the results from Ref. [25], it is consistent with our analysis of the transport peak.
Results are also shown with and without $\pi N$ scattering. When the $\pi N$ interaction is excluded we see the conductivity rise along the phase transition line. This is due to the temperature dropping as one moves from left to right in figure 4. The dropping temperature lowers the $\pi$ and $K$ densities, reducing the $\pi$’s width and increasing the conductivity. This result is similar to what is seen in QED (elaborate). Alternatively, when the $\pi N$ interaction is included we observe a falling conductivity over temperature, because the raising $\mu_{B}$ increases the nucleon density. In the non-relativistic limit nucleons do not transport charge. Therefore, increasing the nucleon density increases the $\pi$’s width, without introducing more charge carries. The nucleons create a resistive medium that the charge carrying pions travel through. Therefore, the dichotomy between temperature and $\mu_{B}$ drastically changes the medium from a more weakly interacting QED-like pion gas, to a strongly interacting nucleon filled medium.

### v2

@@ 1 Introduction
Two approaches to studying hot and dense nuclear matter include the observation and description of electromagnetic probes, and the characterization of the medium with transport coefficients. Dileptons and photons are produced throughout the evolution of the nuclear fireball with little final state interaction, providing a probe of the electromagnetic spectral function ($\rho_{EM}$) in the medium [3]. The production of low-mass dilepton and photon radiation has been well described at multiple experiments [4, 5, 6]. Transport coefficients detail the long-wavelength properties of the medium, describing the flow of conserved charges throughout the collision. In this paper, we calculate the electric conductivity ($\sigma_{el}$), which can be extracted from the zero-energy limit of $\rho_{EM}$ at vanishing 3-momentum [7, 8, 9]. A close relationship exist between electromagnetic probes and the conductivity, as both are proportional to $\rho_{EM}$. In fact, future plans to measure very-low mass and momentum dileptons at the Schwer-Ionen Synchrotron (SIS), RHIC, and the LHC have potential to probe the conductivity’s transport peak [10, 11, 12]. For example, recent work combining hadronic many-body theory with nonperturbative QGP found that a broadening transport peak will induce an enhancement in low-mass dilepton spectra that may be observable in future experiments at the LHC [13]. These initiatives encourage the study of the conductivity, especially considering the significant variation in calculations using different methodologies [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@ 4 Gauge invariance in medium
Within the VDM, the $\rho$ meson couples to a conserved current, requiring it to be 4-dimensionally transverse, $q_{\mu}\Sigma_{\rho}^{\mu\nu}=0$. The Ward identities (Eq. 17) must be satisfied by the $\rho\pi\pi$ ($\Gamma_{\mu\,ab3}^{(3)}$) and $\rho\rho\pi\pi$ ($\Gamma_{\mu\nu\,ab33}^{(4)}$) vertices to ensure transversality:
@@
We start by assuming the corrections can be written as a sum of terms proportional to $\Sigma_{\pi}$. For the spacial components of the 3-point corrections, we additionally assume each term is proportional to a momentum $k_{i}$ or $q_{i}+k_{i}$, and can be written as a function of $k$ or $q+k$, as was the case in Ref. [27]. For the 3-point vertex we arrive at the corrections:
@@
where, $R=\rho,\sigma,K$, and we have suppressed the 3 in the isospin indices for simplicity. Here, we have divided each term in Eq. 22 by either $k_{0}|\vec{k}|$ or $(q_{0}+k_{0})|\vec{q}+\vec{k}|$, so that the corrections will have units of energy, and each term will be a function of either $k$ or $q+k$. This choice assures the corrections are finite in the limit where $k_{0}$ or $|\vec{k}|$ approach zero, without introducing additional parameters. These restrictions combined with Eq. 19 determine effective 3-point corrections. For the 4-point corrections we assume $\Gamma_{ij\,ab}^{\prime(4\,R\pi)}=0$, because its contribution to the transport peak was found to be small in Ref. [26]. This assumption significantly simplifies the 4-point corrections, and allows us to use Eq. 20 to fix $\Gamma_{i0\,ab}^{\prime(4\,\pi)}$ and $\Gamma_{00\,ab}^{\prime(4\,\pi)}$, giving:
@@ 5 EM spectral function at finite temperature and density
Finally, we extract the conductivity using Eq. 9. In the right panel of figure 4 we plot the conductivity over temperature along the phase transition line. We provide results with and without vertex corrections, which provide a $\sim 15\%$ increase. The increase in the conductivity is due to the corrections introducing additional channels that the EM current can couple to the medium through [26]. Although, this result is somewhat larger than the results from Ref. [26], it is consistent with our analysis of the transport peak.
Results are also shown with and without $\pi N$ scattering. When the $\pi N$ interaction is excluded we see the conductivity rise along the phase transition line. This is due to the temperature dropping as one moves from left to right in figure 4. The dropping temperature lowers the $\pi$ and $K$ densities, reducing the $\pi$’s width and increasing the conductivity. Alternatively, when the $\pi N$ interaction is included we observe a falling conductivity over temperature, because the raising $\mu_{B}$ increases the nucleon density. In the non-relativistic limit nucleons do not transport charge. Therefore, increasing the nucleon density increases the $\pi$’s width, without introducing more charge carries. The nucleons create a resistive medium that the charge carrying pions travel through. Therefore, the dichotomy between temperature and $\mu_{B}$ drastically changes the medium from a more weakly interacting pion gas, to a strongly interacting nucleon filled medium.
@@ 6 Summary and future work
Acknowledgements
J.A. was partially supported by the U.S. National Science foundation under grants PHY-1913286 and PHY-2209335.
F.G. is supported in part by the U.S. Department of Energy Office of Science under grant No. DE-SC0005131.